# Asymptotics for conformal inference

### v1

@@ 5 Proofs
In this section, we prove the main results of the paper. They rely on a particular decomposition of the FCP/FDP into two processes that further jointly converge. Applying the functional delta method (van der Vaart,, 1998) then allows to conclude.
@@ 5.1.1 $\mathrm{FCP}$ decomposition
To emphasize that the FCP is the e.c.d.f. of $({{p_{i}^{(n)}}})_{i\in\llbracket m\rrbracket}$, we let for all $n\geqslant 1$ and $m\geqslant 1$,
${\widehat{G}}_{m}^{(n)}:=\frac{1}{m}\sum_{i=1}^{m}{\mathds{1}_{{p_{i}^{(n)}}\leqslant t}}=\mathrm{FCP}_{m}^{(n)}(t),\ t\in[0,1].$ (29)
@@
corresponding to the e.c.d.f. of the calibration score sample and test score sample, respectively. Note that we have ${p_{i}^{(n)}}=1-{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}(T_{i})$ almost surely (under Assumption 1 to ensure that there are no ties almost surely).
@@
$=\frac{1}{m}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)<T_{i}}}-\left(1-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t)\right)$
$=1-\frac{1}{m}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)\geqslant T_{i}}}-\left(1-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t)\right)$
@@
$\sqrt{\tau_{n,m}}\left({\widehat{G}}_{m}^{(n)}(t)-G(t)\right)=$ $-\sqrt{\frac{\tau_{n,m}}{m}}\sqrt{m}\left({\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)\right)$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\left(F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t)\right).$ (32)
@@ 5.1.2 Joint convergence
Thanks to (32), in order to obtain a convergence result for $\sqrt{\tau_{n,m}}\left({\widehat{G}}_{m}^{(n)}-G\right)$, we only need to derive the joint convergence of the two processes delineated in the decomposition (32).
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ G\\ G^{\prime}\mathbb{V}\end{pmatrix}\text{ on $\left[D(0,1)\right]^{2}$},$ (33)
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right]\\ \frac{\tau_{n,m}}{{m}}\\ \frac{\tau_{n,m}}{{n}}\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ G\\ G^{\prime}\mathbb{V}\\ \sigma^{2}\\ 1-\sigma^{2}\end{pmatrix}\text{ on $\left[D(0,1)\right]^{2}\times[0,1]^{2}$}.$
@@
Up to consider a subsequence, one can assume that $n/(n+m)\rightarrow\sigma^{2}\in[0,1]$. The result thus follows from Theorem 3.2 in the special case $F_{\tiny\mbox{cal}}=F_{\tiny\mbox{test}}$, because $G=I$ in this case, and $\sigma\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{V}$ is a standard Brownian bridge.
5.1.5 Proof of Theorem 3.3
@@
${\widehat{G}}_{m}^{w,(n)}(t)$ $:=\frac{1}{m}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{{p_{i}^{w,(n)}}\leqslant t}}=\mathrm{FCP}_{m}^{w,(n)}(t),\ t\in\mathbb{R};$ (34)
${\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}(t)$ $:=\frac{\sum_{k=1}^{n}w(S_{k}){\mathds{1}_{S_{k}\leqslant t}}}{w(+\infty)+\sum_{k=1}^{n}w(S_{k})},\ t\in\mathbb{R},$ (35)
@@
$\sqrt{\tau_{n,m}}\left({\widehat{G}}_{m}^{w,(n)}(t)-G^{w}(t)\right)=$ $-\sqrt{\frac{\tau_{n,m}}{m}}\sqrt{m}\left({\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-t)\right)$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\left(F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ{\left({F^{w}_{{\tiny\mbox{cal}}}}\right)^{-1}}(1-t)\right).$ (36)
Thus, the novelty of (36) with respect to (32) is only the presence of ${({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}$ instead of ${({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}$. By Assumption 4, we can show that the family of function =$\mathcal{F}=\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}\right\}$ is $P_{\tiny\mbox{cal}}$-Donsker and $P_{\tiny\mbox{cal}}$-Glivenko-Cantelli. Since $\mathcal{F}$ is Glivenko-Cantelli, the convergence of the test term happens with the same argument in the unweighted case. Because this class is Donsker, using twice the functional delta method gives us that $(\sqrt{n}[{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}-{({F^{w}_{{\tiny\mbox{cal}}}})^{-1}}])_{n}$ converges in distribution to some known random process on the set $D(0,1)$. This result is stated and then proved in Lemma C.1. This leads to the joint convergence Proposition C.2, which is the analogue of Proposition 5.1 in the weighted case. Theorem 3.3 is thus proved by applying the continuous mapping theorem to the decomposition (36).
@@ 5.2.1 $\mathrm{FDP}$ expression
We introduced, as in the works from Genovese and Wasserman, (2004) and Neuvial, (2008), ${\widehat{{G}}}_{m,{0}}^{(n)}$ and ${\widehat{{G}}}_{m,{1}}^{(n)}$ the two e.c.d.f.’s of conformal $p$-values for non-novelties and novelties, respectively:
@@
$\mathrm{FDP}_{m}^{(n)}(t)$ $=\frac{\pi_{0}(m){\widehat{{G}}}_{m,{0}}^{(n)}(t)}{{\widehat{G}}_{m}^{(n)}(t)\vee m^{-1}},\ t\in(0,1),$
$\mathrm{TDP}_{m}^{(n)}(t)$ $={\widehat{{G}}}_{m,{1}}^{(n)}(t),\ t\in(0,1).$
Following Neuvial, (2008), and since the novelty detection procedure from Bates et al., (2023) is the $\mathrm{BH}$ procedure applied to conformal $p$-values, we can be described it as a thresholding procedure with threshold $\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})$, where the functional $\mathcal{T}^{\mathrm{BH}_{\alpha}}$ is defined by
$\mathcal{T}^{\mathrm{BH}_{\alpha}}(F)\coloneqq\sup\left\{t\in[0,1],F(t)\geqslant\frac{t}{\alpha}\right\}.$ (40)

### v3

@@ 5 Proofs
In this section, we prove the main results of the paper. They rely on a particular decomposition of the FCP/FDP into two processes that further jointly converge. Applying the functional delta method (van der Vaart, 1998) then allows to conclude.
@@ 5.1.1 $\mathrm{FCP}$ decomposition
To emphasise that the FCP is the e.c.d.f. of $({{p_{i}^{(n)}}})_{i\in\llbracket m\rrbracket}$, we let for all $n\geqslant 1$ and $m\geqslant 1$,
${\widehat{G}}_{m}^{(n)}\coloneqq\frac{1}{m}\sum_{i=1}^{m}{\mathds{1}_{{p_{i}^{(n)}}\leqslant t}}=\mathrm{FCP}_{m}^{(n)}(t),\ t\in[0,1].$ (33)
@@
corresponding to the e.c.d.f. of the calibration score sample (with $\infty$) and test score sample, respectively. Note that we have ${p_{i}^{(n)}}=1-{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}(T_{i})$ almost surely (under Assumption 1 to ensure that there are no ties almost surely).
@@
$=\frac{1}{m}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)<T_{i}}}-\mathopen{}\mathclose{{\left(1-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t)}}\right)$
$=1-\frac{1}{m}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)\geqslant T_{i}}}-\mathopen{}\mathclose{{\left(1-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t)}}\right)$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left({\widehat{G}}_{m}^{(n)}(t)-G(t)}}\right)=$ $-\sqrt{\frac{\tau_{n,m}}{m}}\sqrt{m}\mathopen{}\mathclose{{\left({\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)}}\right)$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t)}}\right).$ (36)
@@ 5.1.2 Joint convergence
Thanks to (36), in order to obtain a convergence result for $\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left({\widehat{G}}_{m}^{(n)}-G}}\right)$, we only need to derive the joint convergence of the two processes delineated in the decomposition (36).
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ G\\ G^{\prime}\mathbb{V}\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$},$ (37)
@@
To prove Theorem 3.1, we show the following more general result.
Theorem 5.2.
Assume that for all $n,m$ the calibration and test scores $(S_{1},\dots,S_{n},T_{1},\dots,T_{m})$ are exchangeable without ties almost surely. Then
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{(n)}-I}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\text{ on $D[0,1]$,}$
The two assumptions in Theorem 5.2 are implied by Assumption 1 with $F_{\tiny\mbox{cal}}=F_{\tiny\mbox{test}}$, hence Theorem 3.1 is implied by Theorem 5.2. Let us prove the latter. Since the calibration and test scores are exchangeable without ties a.s. by Proposition 2.2 from Gazin, Blanchard and Roquain (2024) we get for all $n,m\in\mathbb{N}^{2}$ that the conformal $p$-values $({p_{i}^{(n)}})_{i\in\llbracket m\rrbracket}$ has a distribution only depending on $n$ and $m$, denoted by $P_{n,m}$. Hence, for all $T:D[0,1]\rightarrow\mathbb{R}$ bounded continuous funtion we have :
$\mathbb{E}\mathopen{}\mathclose{{\left[T\mathopen{}\mathclose{{\left(\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{(n)}-I}}\right)}}\right)}}\right]=\mathbb{E}\mathopen{}\mathclose{{\left[T\mathopen{}\mathclose{{\left(\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\widetilde{\mathrm{FCP}}_{m}^{(n)}-I}}\right)}}\right)}}\right],$ (38)
where $\widetilde{\mathrm{FCP}}_{m}^{(n)}$ is the $\mathrm{FCP}$ obtained with i.i.d scores uniformly distributed on $(0,1)$. Those specific scores satisfy Assumptions 1 and 3 with $F_{\tiny\mbox{cal}}=F_{\tiny\mbox{test}}=I$ on $[0,1]$, and by (38) we only have to prove the convergence in this specific setting.
Up to consider a subsequence, one can assume that $n/(n+m)\rightarrow\sigma^{2}\in[0,1]$ by compacity of $[0,1]$. By applying Proposition 5.1 with the i.i.d. uniformly distributed on $(0,1)$ scores from the right hand side of (38), we get by Slutsky’s Lemma:
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I)}}\right]\\ \frac{\tau_{n,m}}{{m}}\\ \frac{\tau_{n,m}}{{n}}\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\\ \mathbb{V}\\ \sigma^{2}\\ 1-\sigma^{2}\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}\times[0,1]^{2}$},$
with $(\mathbb{U},\mathbb{V})$ two independent standard Brownian bridges. Applying the continuous mapping theorem with the decomposition (36) gives the following convergence:
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\widetilde{\mathrm{FCP}}_{m}^{(n)}-I}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{V}\overset{\mathcal{L}}{=}\mathbb{W},$
with $\mathbb{W}$ a standard Brownian bridge. By universality of the limiting distribution and compacity of $[0,1]$, the convergence in distribution holds without considering any subsequence.
5.1.4 Proof of Theorem 3.2
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right]\\ \frac{\tau_{n,m}}{{m}}\\ \frac{\tau_{n,m}}{{n}}\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ G\\ G^{\prime}\mathbb{V}\\ \sigma^{2}\\ 1-\sigma^{2}\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}\times[0,1]^{2}$}.$
@@
${\widehat{G}}_{m}^{w,(n)}(t)$ $\coloneqq\frac{1}{m}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{{p_{i}^{w,(n)}}\leqslant t}}=\mathrm{FCP}_{m}^{w,(n)}(t),\ t\in\mathbb{R};$ (39)
${\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}(t)$ $\coloneqq\frac{\sum_{k=1}^{n}w(S_{k}){\mathds{1}_{S_{k}\leqslant t}}}{w(+\infty)+\sum_{k=1}^{n}w(S_{k})},\ t\in\mathbb{R},$ (40)
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left({\widehat{G}}_{m}^{w,(n)}(t)-G^{w}(t)}}\right)=$ $-\sqrt{\frac{\tau_{n,m}}{m}}\sqrt{m}\mathopen{}\mathclose{{\left({\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-t)}}\right)$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-t)-F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({F^{w}_{{\tiny\mbox{cal}}}}}}\right)^{-1}}(1-t)}}\right).$ (41)
Thus, the novelty of (41) with respect to (36) is only the presence of $({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}$} instead of $({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}$}. By Assumption 4, we can show that the family of function $\mathcal{F}=\mathopen{}\mathclose{{\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}}}\right\}$ is $P_{\tiny\mbox{cal}}$-Donsker and $P_{\tiny\mbox{cal}}$-Glivenko-Cantelli. Since $\mathcal{F}$ is Glivenko-Cantelli, the convergence of the test term happens with the same argument in the unweighted case. Because this class is Donsker, using twice the functional delta method gives us that $(\sqrt{n}[{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}-{({F^{w}_{{\tiny\mbox{cal}}}})^{-1}}])_{n}$ converges in distribution to some known random process on the set $D(0,1)$. This result is stated and then proved in Lemma C.1. This leads to the joint convergence Proposition C.2, which is the analogue of Proposition 5.1 in the weighted case. Theorem 3.3 is thus proved by applying the continuous mapping theorem to the decomposition (41).
@@ 5.2.1 $\mathrm{FDP}$ expression
We introduced, as in the works from Genovese and Wasserman (2004) and Neuvial (2008), ${\widehat{{G}}}_{m,{0}}^{(n)}$ and ${\widehat{{G}}}_{m,{1}}^{(n)}$ the two e.c.d.f.’s of conformal $p$-values for non-novelties and novelties, respectively:
@@
$\begin{split}\mathrm{FDP}_{m}^{(n)}(t)&=\frac{\pi_{0}(m){\widehat{{G}}}_{m,{0}}^{(n)}(t)}{{\widehat{G}}_{m}^{(n)}(t)\vee m^{-1}},\ t\in(0,1),\\ \mathrm{TDP}_{m}^{(n)}(t)&={\widehat{{G}}}_{m,{1}}^{(n)}(t),\ t\in(0,1).\end{split}$ (45)
Following Neuvial (2008), and since the novelty detection procedure from Bates et al. (2023) is the $\mathrm{BH}$ procedure applied to conformal $p$-values, we can be described it as a thresholding procedure with threshold $\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})$, where the functional $\mathcal{T}^{\mathrm{BH}_{\alpha}}$ is defined by
$\mathcal{T}^{\mathrm{BH}_{\alpha}}(F)\coloneqq\sup\mathopen{}\mathclose{{\left\{t\in[0,1],F(t)\geqslant\frac{t}{\alpha}}}\right\}.$ (46)